# What should a neuron aim for? Designing local objective functions based on information theory

### v3

@@ 4 Experiments
We now explain how the general concept of trivariate infomorphic neurons can be practically utilized to construct networks solving a supervised classification task, i.e., the MNIST supervised handwritten image classification task. To this end, we demonstrate how an infomorphic network with one hidden layer can be set up using either heuristic or optimized goal function parameters.
Neuron structure and activation function
@@
To construct infomorphic networks, a number of practical decisions need to be made for how the output signal is constructed from the inputs in the forward pass. As a first step, the higher-dimensional signals from the different input classes are aggregated to the single numbers $F$, $C$ and $L$ by a weighted sum. Subsequently, these aggregated inputs are passed into an activation function, which can be chosen arbitrarily, but needs to fit the requirements of the specific application. For the supervised classification task at hand, the function $A(F,C,L)=F[(1-\alpha_{1}-\alpha_{2})+\alpha_{1}\,\sigma(\beta_{1}FC)+\alpha_{2}\,\sigma(\beta_{2}FL)]$ has been chosen, where $\sigma$ refers to the sigmoid function and $\alpha_{1}=\alpha_{2}=0.1$ and $\beta_{1}=\beta_{2}=2$ are parameters that shape the influence of the input compartments on the activation function. This activation function builds on concepts of [18] and makes a clear distinction between the driving feedforward input and the modulatory context and lateral inputs, ensuring that the network performs similarly during training, when the context inputs is provided, and for evaluation, where it is withheld (i.e., by setting $C=0$). To show the independence of the infomorphic approach from a particular activation function, the results are additionally compared to a simple linear activation function $A(F,C,L)=F+0.1L+0.1C$. The output of the activation function is finally mapped to the unit interval by a sigmoid function, whose output is used as a probability to stochastically output “1” (firing) or “-1” (non-firing). For the bivariate output layer, the sigmoid is not sampled from but just interpreted as a firing probability, with the neuron with the highest value being used as the network prediction. The lateral connections between neurons add recurrence to the network, requiring for the same input to be shown multiple times to ensure convergence of the activations. In practice, however, we have seen that presenting the same input twice is sufficient to achieve proper network function (see Section A.9). The inference procedure for infomorphic networks is summarized in pseudocode in Section A.1.
Discretization and local gradients
@@
To optimize a PID-based goal function during training, the first step is to estimate the PID atoms. Since the original $I^{\mathrm{sx}}_{\cap}$ measure only works with discrete variables, the aggregated inputs from a batch of samples are first discretized to $20$ levels each (see Section A.11). Subsequently, the empirical probability mass function is constructed from the batch, from which the PID atoms are computed by means of a plug-in estimation. Finally, the current value of the goal function can determined as a linear combination of PID atoms (Algorithm 2 and Algorithm 3 in Section A.1). For the local gradient computation, the gradient of the individual neuron’s objective function is backpropagated via the output to produce the neuron’s weight updates.
Choosing goal parameters
@@
The parameters $\bm{\gamma}$ for a PID goal function can be obtained in two distinct ways - either heuristically from intuitive notions about the nature of the computations necessary to solve the problem, as explained before, or optimized using a suitable hyperparameter optimization procedure. In this work, we contrast both approaches. For the hyperparameter optimization, we compared two established techniques: The tree-structured Parzen Estimator (TPE, 19), that splits the sample points into two groups with high and low performance, fits a model to each group and draws the next sample points according to the ratio of probabilities, and the Covariance Matrix Adaptation Evolution Strategy (CMA-ES, 20), which generates new evaluation points by sampling from a multivariate Gaussian distribution that was fitted to the best samples from previous iterations.
Network Setup
To solve the MNIST task [21], we devised a network architecture consisting of an input layer, a hidden layer made up from $N_{\mathrm{hid}}$ trivariate infomorphic neurons and an output layer consisting of $10$ bivariate infomorphic neurons. The trivariate neurons in the hidden layer each receive the whole image as their feedforward signal and the output of other neurons of the same layer through lateral connections. The bivariate neurons in the output layer receive all outputs of the hidden layer and additionally exactly one element of the one-hot label as context, which is only provided during training, and are trained to maximize redundancy between their input and the information about the one label they observe.
@@
We explored three slightly different approaches to how the hidden layer can be set up using trivariate infomorphic neurons, illustrated in Fig. 3.A. In setup 1, we use the fully connected label as $C$ and the output of all neurons of the same layer as all-to-all connected lateral input $L$. In setup 2, a sparse connectivity of a maximum of 100 connections is used instead of all-to-all for $L$. In setup 3, additionally to sparse $L$, the label is replaced by a fully connected feedback from the output layer as context $C$, indicating a path towards how multiple hidden layers may be stacked in the future. For comparison, we trained a benchmark network with the same connectivity as setup 1 using standard backpropagation and cross-entropy loss. Additionally, we trained the output layer of a fixed random hidden layer setup using step-function activation.
@@ 5 Results
The performance of all major setups for $N_{\mathrm{hid}}=100$ is shown in Fig. 3.B. The three trivariate infomorphic setups use the same set of optimized goal parameters (see below) for all hidden layer sizes and significantly outperform the random baseline as well as the networks with a bivariate setup for the hidden layer. As shown in Fig. 3.C, Setup 1 matches the performance of backpropagation for up to 100 neurons, and reaches its maximum test accuracy for 500 neurons before its performance starts decreasing with larger layer sizes. This decrease in accuracy can be attributed to a lack of convergence of the neurons (appendix Fig. 5), likely arising through the interaction of too many neurons. To alleviate the convergence problem, we performed additional experiments (setup 2) with the number of lateral connections reduced to maximally 100, which leads to better convergence and contributes to the continuous increase of performance for larger hidden layers, reaching a median test accuracy of $97.5\%$ for 2000 neurons, slightly below the $98.0\%$ of the corresponding backpropagation benchmark. Finally, experiments with setup 3 provide evidence that the direct label input to the hidden layer can be replaced with feedback from the next layer, while still enabling solid performance especially for large layers. Additionally, we tested the alternate linear activation function as outlined in 4. Here, experiments using the optimized goal function parameters resulted in accuracies between $94.9\,\%$ and $95.1\,\%$ for $100$ neurons, which is on par with the original activation function.
@@
Figure 4: Given a set of optimized goal parameters, an ablation study allows for the identification and interpretation of the most critical neural subgoals for a given task. A. The heuristically defined goal function (see Fig. 2.D) shows 92.8% test accuracy in MNIST for $N_{hid}=100$; optimizing the goal function using hyperparameter optimization increases the test accuracy to 94.9%. The optimized goal parameters include the heuristically found $\gamma_{\{F\}\{C\}}$, but also additional PID atoms to be maximized or minimized at the same time. B. For identifying the most important goal parameters, we performed an ablation study (individually setting $\gamma$ parameters to 0) and measured the change in validation accuracy compared to a network trained with the full goal. C. A successive ablation of parameters in order from lowest to highest individual effect identifies four parameters as being crucial for network performance (see Table 3 for their definition and interpretation). The lines indicate mean values, with the intervals depicting the maximum and minimum of 10 runs. D. To test whether more complex image classification tasks require different atoms to be maximized, we perform a separate hyperparameter optimization for setup 1, 100 neurons in CIFAR10 and reach a median test accuracy of $42.5\%$ (compared to $42.2\%$ using backprop and $41.1\%$ using the goal function originally optimized for MNIST).
Goal parameters
@@
The optimized goal function parameters used for training the main three setups at all sizes were obtained by maximizing the validation set accuracy using setup 1 for $N_{\mathrm{hid}}=100$ neurons. We performed a total of six hyperparameter optimizations of 1000 trials, three using TPE and three using CMA-ES samplers, respectively (see Appendix Fig. 6). The best results were achieved using the parameters shown in Fig. 4. The optimized objective functions outperform the heuristic goal function $G_{\text{heur}}=\Pi_{\{F\}\{C\}}$ by including additional PID atoms besides the intuitively derived $\Pi_{\{F\}\{C\}}$. To compare to a more complex image classification task, a single CMA-ES optimization run has been performed for the CIFAR10 task [22], with the optimal parameters presented in Fig. 4.D.
Parameter importance
@@
An analysis of the optimized hyperparameter values (Fig. 4.B) shows that only relatively few parameters are of high importance for performance. By setting goal parameters to zero individually (Figure 4B, right) we find that there are four goal parameters that are critically non-zero for high performance on MNIST. This finding is confirmed by cumulatively setting goal function parameters to zero in the order of the individual drop in performance, which is illustrated in Fig. 4.C. Interestingly, we find that CIFAR10 two of the four critical goal parameters for MNIST seem especially important. As a complementary measure of parameter importance, the results of mean decrease in impurity [23] can be found in Section A.6. The results of a more detailed analysis of parameter importance, including the effects of varying parameter values, are presented in Fig. 9.
Interpretation
@@
For the MNIST task, neurons aim for (i) encoding information that is redundant between feedforward and context inputs and thus task relevant, but not already encoded in other neurons (strongly positive $\gamma_{\{F\}\{C\}}$), (ii) avoid encoding information that is not in the context, thus not task-relevant, and already encoded by other neurons (strongly negative $\gamma_{\{F\}\{L\}}$) and (iii) allow redundancies of task relevant information between neurons (slightly positive $\gamma_{\{F\}\{C\}\{L\}}$). Additionally, neurons avoid synergies that require $F$ and either $C$ or $L$ for their recovery ($\gamma_{\{FC\}\{FL\}}$) (see Table 3 for an overview of the parameter values and interpretations). For the CIFAR10 task (see Fig. 4.D), the two most important parameters $\gamma_{\{F\}\{C\}\{L\}}$ and $\gamma_{\{F\}\{C\}}$ point to a prioritization of redundancy over uniqueness between neurons.
@@ 6 Related Works
We are not aware of any other attempts to use PID directly as a constructive tool for the design of ANNs with local learning objectives, other than the bivariate implementation of infomorphic networks [17]. The most directly related work using other local information-theoretic learning objectives is possibly the work on coherent infomax by Kay, Phillips and colleagues [24]. Their idea of defining a neural goal function that ensures only information coherent between the inputs being passed on in the output of a neuron inspired our idea of using PID-based redundancy as a neural goal function. However, as analyzed in [7], the lack of a PID at the time of invention of coherent infomax reduced the expressiveness of their goal function, and also prevented an easily interpretable formulation for more than two input classes. As shown here, however, three input classes seem to be necessary to unlock the potential of information theoretic learning, in the sense that a neuron needs to know at least the data to transform ($F$), the relevance of various aspects of these data ($C$), and also what is not yet encoded well by other neurons (via $L$).
@@
Aside from PID proper, there are two prominent principles based on information theory that has been employed to train deep neural networks: information maximization (InfoMax, 25) and information bottleneck (IB, 26). The main idea of InfoMax is to learn representations that encode as much information as possible about the task, while IB additionally penalizes information that is irrelevant to the task [27]. Various frameworks have employed InfoMax and IB where unlike in our framework, information theory is employed at the level of group of neurons (layer). Examples include Deep InfoMax where mutual information of high-level representations and the input is maximized [28], InfoNCE that maximizes the mutual information of future observations and the current latent representation [29], and Variational autoencoders that employ an IB layer [30].
@@
Additionally, local learning rules based on concepts other than information theory have been developed and show potential for scaling to larger, more capable networks [31, 32, 33, 34]. This includes learning rules based on concepts from contrastive learning [35, 36], predictive coding [37, 38, 39], and many others [40, 41, 42, 43, 44, 45]. Most of these approaches are confined to specific learning paradigms and implementations, while our framework describes local learning goals in an abstract manner, making it applicable to various learning paradigms, datasets and implementations, while still being interpretable.
@@
In terms of abstract per-neuron objectives, recently very interesting complementary work has been published, building on concepts from control theory [46] and reinforcement learning  [47]. These efforts are in line with our work, and we believe similar creative, yet principled approaches could provide a basis for deepening the understanding of local information processing.
@@
Beyond the formulation of information theoretic learning rules, there is literature using PID to analyze deep neural network function, such as [16, 15, 48, 49, 50, 51]. These studies necessarily differ from the perspective on neural function taken here, as they analyze the function of neurons in feedforward networks trained with backprop, where – different to the infomorphic approach – the learning signals are not directly available as inputs to the neural activation function.
@@ 7 Discussion
In this work we demonstrate how neurons can self-organize to solve classification tasks using local PID-based goal functions. These local information-theoretic objectives of neural learning are formulated in human interpretable terms – in contrast to the usual, more mechanistic description of local learning rules, or the formulation of a purely global objective function. Our framework expresses this local objective by selecting which output information should be uniquely, redundantly or synergistically determined by the various classes of local inputs to a neuron. We show how to derive the learning objective for supervised learning based on intuitive heuristic reasoning, and upon some optimization of this simple rule demonstrate performance on MNIST that is comparable to the same ANN trained with backpropagation.
@@
Limitations and Outlook
We show that the infomorphic local optimization approach reaches a performance comparable to backpropagation in a setup with one fully connected hidden layer. In subsequent work, we will investigate the local information processing and connectivity required for deeper networks. Preliminary experiments show that deeper networks train out of the box using the same goal function parameters and reach the same or slightly higher accuracy compared to the single hidden layer models (see subsection A.10). More work is needed to determine how this accuracy can be improved, for instance by choosing different goal function parameters for the different layers.
The infomorphic neurons presented in this work use a discrete version of PID, which limits the amount of information that individual neurons can convey. While this is similar to the binary output states of “spike” and “no spike” in biological neurons, PID can also be applied to continuous variables and thus could also enable learning in continuous neurons [52].
For the study of biological neural networks, infomorphic networks could potentially serve as a powerful constructive modeling approach. While information-theoretic gradients are almost certainly too complex to be directly implemented in a biological substrate, biological neurons may still strive for similar objectives through simpler approximations and mechanistic local learning rules. Having a principled and general way to formulate, identify and test local information processing objectives could provide a new perspective on local learning in general. Additionally, subsequent work could aim to establish a formal bridge between local goals and known local learning rules, reconciling well-established Hebbian-like rules with novel principled approaches.
@@
Code Availability
Acknowledgments and Disclosure of Funding
We would like to thank Marcel Graetz, Jonas Dehning, Mark Blümel, Fabian Mikulasch and Lucas Rudelt for their input and fruitful discussions about this topic. We would also like to thank Johannes Zierenberg and Sebastian Mohr and the rest of the Priesemann and the Wibral Group for their valuable comments and feedback on this work. A.S., V.N. and V.P. were funded via the MBExC by the Deutsche Forschungsgemeinschaft (DFG) under Germany’s Excellence Strategy-EXC 2067/1-390729940. A.S., V. N., V.P., and M.W., were supported and funded by the DFG via the RTG 2906 “Curiosity” project ID 502807174. D.E. and M.W. were supported by a funding from the Ministry for Science and Education of Lower Saxony and the Volkswagen Foundation through the “Niedersächsisches Vorab” under the program “Big Data in den Lebenswissenschaften” – project “Deep learning techniques for association studies of transcriptome and systems dynamics in tissue morphogenesis”. A.M. and M.W. are employed at the Campus Institute for Dynamics of Biological Networks (CIDBN) funded by the Volkswagenstiftung. A.E., V.P. and M.W. received funding from the DFG via the SFB 1528 “Cognition of Interaction” - project-ID 454648639. A.E. was supported from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation programme (Grant agreement No. 101041669). M.W. was supported by the flagship science initiative of the European Commission’s Future and Emerging Technologies program under the Human Brain project, HBP-SP3.1-SGA1-T3.6.1.

### v4

@@ 4 Experiments
We now explain how the general concept of trivariate infomorphic neurons can be practically utilized to construct networks solving a supervised classification task, i.e., the MNIST supervised handwritten digit classification task. To this end, we demonstrate how an infomorphic network with one hidden layer can be set up using either heuristic or optimized goal function parameters.
@@
Neuron structure and activation function   To construct infomorphic networks, a number of practical decisions need to be made for how the output signal is constructed from the inputs in the forward pass. As a first step, the higher-dimensional signals from the different input classes are aggregated to the single numbers $F$, $C$ and $L$ by a weighted sum. Subsequently, these aggregated inputs are passed into an activation function, which can be chosen arbitrarily, but needs to fit the requirements of the specific application. For the supervised classification task at hand, the function $A(F,C,L)=F[(1-\alpha_{1}-\alpha_{2})+\alpha_{1}\,\sigma(\beta_{1}FC)+\alpha_{2}\,\sigma(\beta_{2}FL)]$ has been chosen, where $\sigma$ refers to the sigmoid function and $\alpha_{1}=\alpha_{2}=0.1$ and $\beta_{1}=\beta_{2}=2$ are parameters that shape the influence of the input compartments on the activation function. This activation function builds on concepts of Kay (1994) and makes a clear distinction between the driving feedforward input and the modulatory context and lateral inputs, ensuring that the network performs similarly during training, when the context inputs is provided, and for evaluation, where it is withheld (i.e., by setting $C=0$). To show the independence of the infomorphic approach from a particular activation function, the results are additionally compared to a simple linear activation function $A(F,C,L)=F+0.1L+0.1C$. The output of the activation function is finally mapped to the unit interval by a sigmoid function, whose output is used as a probability to stochastically output “1” (firing) or “-1” (non-firing). For the bivariate output layer, the sigmoid is not sampled from but just interpreted as a firing probability, with the neuron with the highest value being used as the network prediction. The lateral connections between neurons add recurrence to the network, requiring for the same input to be shown multiple times to ensure convergence of the activations. In practice, however, we have seen that presenting the same input twice is sufficient to achieve proper network function (see Section A.9). The inference procedure for infomorphic networks is summarized in pseudocode in Section A.1.
@@
Discretization and local gradients   To optimize a PID-based goal function during training, the first step is to estimate the PID atoms. Since the original $I^{\mathrm{sx}}_{\cap}$ measure only works with discrete variables, the aggregated inputs from a batch of samples are first discretized to $20$ levels each (see Section A.11). Subsequently, the empirical probability mass function is constructed from the batch, from which the PID atoms are computed by means of a plug-in estimation. Finally, the current value of the goal function can determined as a linear combination of PID atoms (Algorithm 2 and Algorithm 3 in Section A.1). For the local gradient computation, the gradient of the individual neuron’s objective function is backpropagated via the output to produce the neuron’s weight updates.
@@
Choosing goal parameters   The parameters $\bm{\gamma}$ for a PID goal function can be obtained in two distinct ways—either heuristically from intuitive notions about the nature of the computations necessary to solve the problem, as explained before, or optimized using a suitable hyperparameter optimization procedure. In this work, we contrast both approaches. For the hyperparameter optimization, we compared two established techniques: The Tree-structured Parzen Estimator (TPE, Bergstra et al., 2011), that splits the sample points into two groups with high and low performance, fits a model to each group and draws the next sample points according to the ratio of probabilities, and the Covariance Matrix Adaptation Evolution Strategy (CMA-ES, Hansen, 2016), which generates new evaluation points by sampling from a multivariate Gaussian distribution that was fitted to the best samples from previous iterations.
@@
Network Setup   To solve the MNIST task (LeCun et al., 1998), we devised a network architecture consisting of an input layer, a hidden layer made up from $N_{\mathrm{hid}}$ trivariate infomorphic neurons and an output layer consisting of $10$ bivariate infomorphic neurons. The trivariate neurons in the hidden layer each receive the whole image as their feedforward signal, a context signal that guides their learning, and the output of other neurons of the same layer through lateral connections. The bivariate neurons in the output layer receive all outputs of the hidden layer and additionally exactly one element of the one-hot label as context, which is only provided during training, and are trained to maximize redundancy between their input and the information about the one label they observe. We explored three slightly different approaches to how the hidden layer can be set up using trivariate infomorphic neurons, illustrated in Fig. 3.A. In setup 1, we use the fully connected one-hot label as $C$ and the output of all neurons of the same layer as all-to-all connected lateral input $L$. In setup 2, a sparse connectivity of a maximum of 100 connections is used instead of all-to-all for $L$. In setup 3, additionally to sparse $L$, the label is replaced by a fully connected feedback from the output layer as context $C$, indicating a path towards how multiple hidden layers may be stacked in the future. For comparison, we trained a benchmark network with the same connectivity as setup 1 using standard backpropagation and cross-entropy loss. Additionally, we trained the output layer of a fixed random hidden layer setup using step-function activation.
@@ 5 Results
Performance   The performance of all major setups for $N_{\mathrm{hid}}=100$ is shown in Fig. 3.B. The three trivariate infomorphic setups use the same set of optimized goal parameters (see below) for all hidden layer sizes and significantly outperform the random baseline as well as the networks with a bivariate setup for the hidden layer. As shown in Appendix Fig. 3.C, Setup 1 matches the performance of backpropagation for up to 100 neurons, and reaches its maximum test accuracy for 500 neurons before its performance starts decreasing with larger layer sizes. This decrease in accuracy can be attributed to a lack of convergence of the neurons (Appendix Fig. 5), likely arising through the interaction of too many neurons. To alleviate the convergence problem, we performed additional experiments (setup 2) with the number of lateral connections reduced to maximally 100, which leads to better convergence and contributes to the continuous increase of performance for larger hidden layers, reaching a median test accuracy of $97.5\%$ for 2000 neurons, slightly below the $98.0\%$ reached via backpropagation. Finally, experiments with setup 3 provide evidence that the direct label input to the hidden layer can be replaced with feedback from the next layer, while still enabling solid performance especially for large layers. Additionally, we tested the alternate linear activation function as outlined in 4. Here, experiments using the optimized goal function parameters resulted in accuracies between $94.9\,\%$ and $95.1\,\%$ for $100$ neurons, which is on par with the original activation function.
@@
Figure 4: Given a set of optimized goal parameters, an ablation study allows for the identification and interpretation of the most critical neural subgoals for a given task. A. The heuristically defined goal function (see Fig. 2.D) shows 92.8% test accuracy on MNIST for $N_{hid}=100$; optimizing the goal function using hyperparameter optimization increases the test accuracy to 94.9%. The optimized goal parameters include the intuitive $\gamma_{\{F\}\{C\}}$, but also additional PID atoms to be maximized or minimized at the same time. B. For identifying the most important goal parameters, we performed an ablation study (individually setting $\gamma$ parameters to 0) and measured the change in validation accuracy compared to a network trained with the full goal. C. A successive ablation of parameters in order from lowest to highest individual effect identifies four parameters as being crucial for network performance (see Table 3 for their definition). The lines indicate mean values, with the intervals depicting the maximum and minimum of 10 runs. D. To test whether more complex image classification tasks require different goal parameters, we perform a separate hyperparameter optimization for setup 1, 100 neurons in CIFAR-10 and reach a median test accuracy of $42.5\%$ (compared to $42.2\%$ using backprop and $41.1\%$ using the goal function optimized for MNIST).
@@
Goal parameters   The optimized goal function used for training the three main setups at all sizes were obtained by maximizing the validation set accuracy using setup 1 with $N_{\mathrm{hid}}=100$. We performed a total of six hyperparameter optimizations of 1000 trials, three using TPE and three using CMA-ES samplers, respectively (see Appendix Fig. 6). The best results were achieved using the parameters shown in Fig. 4. The optimized objective functions outperform the intuitive goal function by including additional PID atoms besides the intuitively derived $\Pi_{\{F\}\{C\}}$. To compare to a more complex image classification task, a single CMA-ES optimization run has been performed for the CIFAR-10 task (Krizhevsky, 2012), with the optimal parameters presented in Fig. 4.D.
@@
Parameter importance   An analysis of the optimized hyperparameter values (Fig. 4.B) shows that only relatively few parameters are of high importance for performance. By setting goal parameters to zero individually (Figure 4.B) we find that there are four goal parameters that are critically non-zero for high performance on MNIST. This finding is confirmed by cumulatively setting goal function parameters to zero in the order of the individual drop in performance, which is illustrated in Fig. 4.C. Interestingly, we find that CIFAR-10 two of the four critical goal parameters for MNIST seem especially important. As a complementary measure of parameter importance, the results of mean decrease in impurity (Breiman, 2001) can be found in Section A.6. The results of a more detailed analysis of parameter importance are presented in Appendix Fig. 9.
@@
Interpretation   For the MNIST task, neurons aim for (i) encoding information that is redundant between feedforward and context inputs and thus task relevant, but not already encoded in other neurons (strongly positive $\gamma_{\{F\}\{C\}}$), (ii) avoid encoding information that is not in the context, thus not task-relevant, and already encoded by other neurons (strongly negative $\gamma_{\{F\}\{L\}}$) and (iii) allow redundancies of task relevant information between neurons (slightly positive $\gamma_{\{F\}\{C\}\{L\}}$). Additionally, neurons avoid synergies that require $F$ and either $C$ or $L$ for their recovery ($\gamma_{\{FC\}\{FL\}}$) (see Appendix Table 3 for an overview of the parameter values and interpretations). For the CIFAR-10 task (see Fig. 4.D), the two most important parameters $\gamma_{\{F\}\{C\}\{L\}}$ and $\gamma_{\{F\}\{C\}}$ point to a prioritization of redundancy over uniqueness between neurons.
@@ 6 Related Works
We are not aware of any other attempts to use PID directly as a constructive tool for designing ANNs with local learning objectives, aside from the bivariate implementation of infomorphic networks (Makkeh et al., 2025). The most directly related work using other local information-theoretic learning objectives is likely the work on coherent infomax by Kay, Phillips and colleagues (Kay and Phillips, 1997). Their idea of defining a neural goal function that ensures only information coherent between the inputs being passed on in the output of a neuron inspired our idea of using PID-based redundancy as a neural goal function. However, as analyzed in Wibral et al. (2017), the lack of a PID at the time of invention of coherent infomax reduced the expressiveness of their goal function, and also prevented an easily interpretable formulation for more than two input classes. Our results suggest that three input classes are crucial to unlock the potential of information theoretic learning, in the sense that a neuron needs to know at least the data to transform ($F$), the relevance of various aspects of these data ($C$), and also what is not yet encoded well by other neurons (via $L$).
@@
Aside from PID proper, there are two prominent principles based on information theory that has been employed to train deep neural networks: information maximization (InfoMax, Linsker, 1988) and information bottleneck (IB, Tishby et al., 2000). The main idea of InfoMax is to learn representations that encode as much information as possible about the task, while IB additionally penalizes information that is irrelevant to the task (Yu et al., 2020). Various frameworks have employed InfoMax and IB where unlike in our framework, information theory is employed at the level of group of neurons (layer). Examples include Deep InfoMax where mutual information of high-level representations and the input is maximized (Hjelm et al., 2019), InfoNCE that maximizes the mutual information of future observations and the current latent representation (Oord et al., 2018), and Variational autoencoders that employ an IB layer (Alemi et al., 2017).
@@
Additionally, local learning rules based on concepts other than information theory have been developed and show potential for scaling to larger, more capable networks (Lillicrap et al., 2020, Richards et al., 2019, Jeon and Kim, 2023, Bredenberg et al., 2024). This includes learning rules based on concepts from contrastive learning (Illing et al., 2021, Ahamed et al., 2023), predictive coding (Mikulasch et al., 2023, Sacramento et al., 2018, Millidge et al., 2022), and many others (Launay et al., 2020, Hinton, 2022, Høier et al., 2023, Lee et al., 2015, Nøkland and Eidnes, 2019, Lässig et al., 2023). Most of these approaches are confined to specific learning paradigms and implementations, while our framework describes local learning goals in an abstract manner, making it applicable to various learning paradigms, datasets and implementations, while still being interpretable.
@@
In terms of abstract per-neuron objectives, recently complementary work has been published, building on concepts from control theory (Moore et al., 2024) and reinforcement learning  (Ott, 2020). These efforts are in line with our work, and we believe similar creative, yet principled approaches could provide a basis for deepening the understanding of local information processing.
@@
Beyond the formulation of information theoretic learning rules, there is literature using PID to analyze deep neural network function, such as Tax et al. (2017), Ehrlich et al. (2023), Tokui and Sato (2021), Liang et al. (2023; 2024), Mohamadi et al. (2023). These studies necessarily differ from the perspective on neural function taken here, as they analyze the function of neurons in feedforward networks trained with backprop, where—different to the infomorphic approach—the learning signals are not directly available as inputs to the neural activation function.
@@ 7 Discussion
In this work we demonstrate how neurons can self-organize to solve classification tasks using local PID-based goal functions. Unlike mechanistic local learning rules or purely global objectives, our framework expresses learning in human interpretable terms. These local objectives are expressed by selecting which output information should be uniquely, redundantly or synergistically determined by the various classes of local inputs to a neuron. We show how to derive a supervised learning objective based on intuitive heuristic reasoning, and upon some optimization of this simple rule demonstrate performance on MNIST that is comparable to the same ANN trained with backpropagation.
@@
Limitations and Outlook   We show that the infomorphic local optimization approach reaches a performance comparable to backpropagation in a setup with one fully connected hidden layer. In subsequent work, we will investigate the local information processing and connectivity required for deeper networks. Preliminary experiments show that deeper networks train out of the box using the same goal function parameters and reach the same or slightly higher accuracy compared to the single hidden layer models (see subsection A.10). More work is needed to determine how this accuracy can be improved, for instance by choosing different goal function parameters for the different layers.
The infomorphic neurons presented in this work use a discrete version of PID, which limits the amount of information that individual neurons can convey. While this is similar to the binary output states of “spike” and “no spike” in biological neurons, PID can also be applied to continuous variables and thus could also enable learning in continuous neurons (Ehrlich et al., 2024).
For the study of biological neural networks, infomorphic networks could potentially serve as a powerful constructive modeling approach. While information-theoretic gradients are almost certainly too complex to be directly implemented in biological substrate, biological neurons may still strive for similar objectives through simpler approximations and mechanistic local learning rules. Having a principled and general way to formulate, identify and test local information processing objectives could provide a new perspective on local learning in general. Additionally, subsequent work could aim to establish a formal bridge between local goals and known local learning rules, reconciling well-established Hebbian-like rules with novel principled approaches.
@@
Code Availability   The framework and the code to reproduce the results of this work are available under https://github.com/Priesemann-Group/Infomorphic_Networks.
Acknowledgments
We would like to thank Marcel Graetz, Jonas Dehning, Mark Blümel, Fabian Mikulasch and Lucas Rudelt for their input and fruitful discussions about this topic. We would also like to thank Johannes Zierenberg, Sebastian Mohr, Darius Burr and the rest of the Priesemann and the Wibral Group for their valuable comments and feedback on this work. A.S., V.N., A.E. and V.P. were funded via the MBExC by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy-EXC 2067/1-390729940. A.S., V.N., A.E., V.P., and M.W., were supported and funded by the DFG – GRK2906 – project number 502807174. V.N. was partly supported by the Else Kröner Fresenius Foundation via the Else Kröner Fresenius Center for Optogenetic Therapies. D.E. and M.W. were supported by a funding from the Ministry for Science and Education of Lower Saxony and the Volkswagen Foundation through the “Niedersächsisches Vorab” under the program “Big Data in den Lebenswissenschaften” – project “Deep learning techniques for association studies of transcriptome and systems dynamics in tissue morphogenesis”. A.M. and M.W. are employed at the Campus Institute for Dynamics of Biological Networks (CIDBN) funded by the Volkswagenstiftung. A.E., V.P. and M.W. received funding from the DFG via the SFB 1528 “Cognition of Interaction” - project-ID 454648639. A.E. was supported from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation programme (Grant agreement No. 101041669). M.W. was supported by the flagship science initiative of the European Commission’s Future and Emerging Technologies program under the Human Brain project, HBP-SP3.1-SGA1-T3.6.1.